Shift report 20030206 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Feb 06 2003 - 00:25:00 EST

  • Next message: Apache: "Shift report 20030206 00:00-08:00"
    ------------------------------------------------------------
    Date: Thu Feb  6 00:25:00 2003, Shift: 16:00-24:00
    Supervisor: kh, with: 
    Log book#14 pages 1 - 1
    
    Shift Summary:
    ------------------------------------------------------------
    Beam studies until 18:00
    Then they started trying to extract the beam.
    There were several problems.  At one point they had a nice beam, but then lost it due to a permit pulled.
    
    Achieved beam about 11:00 with very high background.  I argued with MCR and they tuned as I was on the telephone and managed to bring it to the outer limits of our acceptance.
    ie 50K in TOFW.  The ratio was not too bad `12
    
    Run 7808 in progress; it completes MRS 1050A 45 deg setting
    Next shift will change it immediately (I might even help)
    
    Run plan:
    MRS
    400K 1050B at 45 deg
    400K 400B at 40 deg
    400K 400A at 40 deg
    
    FS
    300K 1/15B at 8 deg
    200K 1/8B at 8 deg
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Thu Feb 06 2003 - 00:29:26 EST